The stock of United Industries has a beta a 2.26 and an expected return of 12.0. The risk-free rate of return is 4 percent. What is the expected return on the market?

options: 7.66% 8.69% 8.24% 8.89% 7.54%

The expected return on JK stock is 14.00 percent while the expected return on the market is 11.00 percent. The beta of JK stock is 1.5. What is the risk-free rate of return?

options: 5.00 percent 3.90 percent 4.90 percent 4.31 percent 5.90 percent
Answer #1

Expected return=risk free rate+beta*(market rate-risk free rate)

a.

12=4+2.26*(market rate-4)

(12-4)=2.26*(market rate-4)

market rate=(12-4)/2.26+4

=7.54%(Approx).

b.

14=Risk free rate+1.5*(11-Risk free rate)

14=Risk free rate+16.5-1.5*Risk free rate

Risk free rate=(16.5-14)/(1.5-1)

=5%